OpenStudy (anonymous):

Need help FAST 1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

Wasnt this already answered?

OpenStudy (anonymous):

I would subtract 8, then divide by 2.

OpenStudy (anonymous):

System of equations would help T- tony B- belinda Tony and Belinda have a combined age of 56 This means that they both add up to be 56 t + b = 56 Belinda is 8 more than twice Tony’s age Tip: when you see 'is' it usually means = b = 8+ 2t b= 8+2t Now there is a system of equations: b=2t t+b=56 Solve for both by using either elimination or substitution

OpenStudy (anonymous):

no they showed me how to do it but i really need someone to show me the work and answer which i know isnt right but i need help with this cause i have a meeting with a teacher in 5 min and i could go over it before they call

OpenStudy (anonymous):

Ok do you know how to solve b = 2t and t + b = 56?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

* sorry, 8 +2t= b Substitute 8 + 2t for the second equation, where b is supposed to be. t + 8 + 2t = 56 Solve for t